# DistTrain: Addressing Model and Data Heterogeneity with Disaggregated Training for Multimodal Large Language Models

### v2

@@ 6 Addressing Data Heterogeneity
Disaggregated training enables preprocessing the training data with negligible runtime overhead. Based on this, we introduce disaggregated data reordering, seamlessly integrated into data preprocessing, to address data heterogeneity without additional overhead. Initially, we present intra-microbatch reordering to eliminate stragglers across DP groups. We then introduce inter-microbatch reordering to minimize pipeline bubbles caused by the straggler microbatches. The combination of the two reordering, i.e., disaggregated data reordering, effectively addresses the data heterogeneity. We emphasize that these two reordering algorithms are integrated into the disaggregated data preprocessing. This ensures that the complex reordering does not interfere with the GPU training process.
6.1 Intra-microbatch Reordering
@@
Insight. To address the first subproblem, intra-microbatch stragglers, we identify the straggler by pinpointing the DP group with the largest training samples. As illustrated in Figure 6, the first DP group becomes a straggler as it contains the two largest training samples. To neutralize this imbalance, we propose reordering the training samples within the global batch by size. Specifically, as depicted in Figure 11, we reorder the training samples into the sequence . This strategy effectively distributes the computational load more evenly and improves the scalability. It eliminates the straggler without breaking the synchronous training semantics [31], as the reordering occurs within an individual global batch.
Intra-microbatch Reordering. Leveraging this insight, we propose intra-microbatch reordering to balance computational load and improve the overall scalability. Formally, the challenge involves minimizing the maximum computation time among DP groups. This problem corresponds to the NP-hard problem, multiway number partitioning [36], which aims to minimize the largest sum. There is no known polynomial-time algorithm capable of determining the optimal solution. Given the substantial batch size in production training, the algorithm employed must be lightweight and efficient. Consequently, we adopt the greedy number partitioning algorithm, which guarantees an approximation ratio of [37]. The detailed algorithm is summarized in Algorithm 2. The function receives the original training samples and DP size . This algorithm first sorts the training samples in ascending order by the sample size (line 3). Then, it loops over the training samples and assigns the sample to the DP group with the current lowest computational load (line 6-8). It then returns the reordered samples (line 9-11). The algorithm has a time complexity of .
6.2 Inter-microbatch Reordering
@@
The second subproblem is the inter-microbatch straggler. As we discussed in §2.3, data heterogeneity leads to varied computation times across microbatches within the modality encoder and generator. The straggler microbatch prolongs the training by creating large pipeline bubbles. In the context of 1F1B pipeline scheme, the overall iteration time is primarily governed by pipeline bubbles and the computation time at the first PP stage of the modality encoder, as illustrated in Figure 12. Let the PP size be and the number of microbatches be ( and in Figure 12). Typically, is larger than to reduce the proportion of time spent in the warm-up and cool-down phases. We abstract an innovative concept, pipeline intervals, at the first PP stage. As shown in Figure 12, these intervals are typically filled with forward pass, except for the last intervals (i.e., , , and ). Straggler microbatches in either the encoder or generator prolong these intervals or increase the unfilled area (i.e., bubble).
@@
Insights. We leverage two insights to solve this problem. The first insight involves minimizing the volume of intervals that are not filled. As shown in Figure 12, the last intervals (i.e., to ) remain unfilled. These intervals become the pipeline bubbles and prolong the training iteration. We observe a positive correlation between the volume of and the size of the microbatch. The size refers to the computation time of the microbatch in modality encoder and generator. For instance, is significantly larger than and since the microbatch is the largest. By strategically reordering the training samples to position the smallest microbatches at the end, we are able to reduce these unfilled intervals (i.e., pipeline bubbles).
@@
The second insight involves minimizing the unfilled area of left intervals. The left intervals (i.e., to ) are filled with the forward pass. As shown in Figure 12, the first interval is filled with by the to forward passes. For subsequent intervals, is filled by the forward pass. By evaluating the volume of , we place the microbatches, whose forward time most closely matches this volume, at the corresponding position, to minimize the unfilled area (i.e., pipeline bubbles).
@@
Inter-microbatch Reordering. Based on the two insights, we design runtime inter-microbatch reordering to minimize the pipeline bubbles. This algorithm is designed for 1F1B pipeline scheme. We will retrofit the algorithm to VPP (i.e., interleaved 1F1B) later. Algorithm 3 summarizes the pseudo code. The function receives the original order of microbatches and the PP size . It stores reordered microbatches in and pending microbatches in (line 2). Early in the process, the smallest microbatch is placed at first to activate all pipeline stages promptly (line 3). Subsequently, it selects the smallest microbatches and places them at the end to minimize unfilled intervals (line 4 and line 12). It then loops over the remaining microbatches (line 5-11). In each loop iteration, it calculates the volume of the interval through the function . For the first interval, it selects microbatches that closely match the interval volume in sum forward time; for others, it selects a single microbatch whose forward time aligns closely with the interval volume. This loop ensures maximal filling of the remaining intervals, which minimizes pipeline bubbles.
@@
The functions and operate with a time complexity of . The function calculates interval volumes using the current order . This calculation is facilitated by a dynamic programming algorithm that utilizes a recursive formula derived from pipeline dependencies. Specifically, each microbatch’s start time depends on two factors: the completion of the preceding microbatch on the same device and the availability of input data from the upstream microbatch. Consequently, the end time of each microbatch is determined by the maximum of these two dependencies plus its own computation time. This dynamic programming algorithm exhibits a complexity of per function invocation. The algorithm has a time complexity of .
@@
Virtual pipeline parallelism (i.e., interleaved 1F1B) also follows the one forward and one backward pipeline scheme to reduce the memory footprint. The fundamental insights of our algorithm apply universally to any 1F1B-based pipeline, including VPP. We adapt the algorithm by computing multiple (i.e., VPP size) intervals and filling them with the corresponding number of forward passes from a single microbatch.
7 System Implementation and Optimization
We implement DistTrain with 6.3K lines of code in Python and C++, and integrate it with Megatron-LM [17], a state-of-the-art training framework for large language models. DistTrain is able to support multimodal LLM training with a variety of modalities on a large-scale GPU cluster (e.g., thousands of GPUs used in §8). DistTrain leverages a distributed file system to store the training data and model checkpoints. It handles failures by automatically recovering the training process from the latest checkpoint.
@@ 7 System Implementation and Optimization
DistTrain manager. DistTrain’s training manager, implemented as a Python script on a dedicated CPU node, formulates the disaggregated model orchestration problem using Disciplined Convex Programming [38]. It employs the CVX solver [39] to efficiently solve this problem within milliseconds. The manager then records the optimal resource allocation and parallelism strategy to a configuration file, which the Kubernetes controller uses to launch the training task.
DistTrain initializer. DistTrain incorporates parallelism unit that manages specific parallelism strategies for each module in multimodal LLMs. Parallelism unit is initialized using PyTorch Distributed [40] library with NCCL as the communication backend, except for the TP communication where we use StepCCL instead. DistTrain shards the model parameters in accordance with the established parallelism strategy.
Parallelism unit. As we discussed in §4.1, the GPU training disaggregation is implemented through a specific module: parallelism unit. When initializing the distributed training, DistTrain first establishes the communication groups within one parallelism unit. Each GPU process has a global rank and a local rank within the unit, which facilitates the distributed initialization. Then, it initializes the communication broker to establish the PP communication between adjacent parallelism units. All communication traffic between parallelism units is routed via the communication broker.
@@
We implement the communication broker by modifying the batched send and receive operations in Megatron-LM to separate operations. This allows flexible communication between multiple upstream and downstream GPU processes. The communication broker adjusts the communication data by concentrating and scattering the data as needed while maintaining the data order. Strategically located on the GPU of the last PP stage in the upstream unit or the first PP stage in the downstream unit, the communication broker avoids additional communication overhead. Besides, the number of communication brokers between two units is determined by the greatest common divisor of the two parallelism units’ DP sizes to maximize the communication bandwidth. Moreover, in the Megatron-LM, the reliance on synchronous communication compels upstream stages to pause until downstream stages fully receive the data. This introduces unnecessary dependencies in the pipeline. To alleviate this, we implement asynchronous send operations that eliminate these superfluous dependencies and redesign the communication topology to prevent potential deadlocks.
Mitigating TP overhead with StepCCL. Tensor Parallelism (TP) is commonly adopted to facilitate training large Transformer-based models with multiple GPUs connected with high bandwidth (e.g., with NVLinks). Specifically, TP divides the linear layers, into smaller sub-modules, which are then distributed across the GPUs. After parallel computation, all GPUs perform collective communication to aggregate data and produce identical results. The TP communication overhead severely degrades overall performance, especially on A800 and H800 GPUs with restricted NVLink bandwidth.
We implement the communication overlap with an in-house collective communication library called StepCCL to reduce the TP overhead. StepCCL is a PyTorch custom plugin that performs cross-GPU collective communication, including allgather, reduce-scatter, and allreduce, which is similar to NCCL. However, NCCL occupies several CUDA Streaming Multiprocessors (SMs) for executing its communication kernel and is known to harm the performance of its concurrent GEMM (matrix multiplication) [41]. To solve this, StepCCL leverages the DMA engine directly to perform data transmission without using any SM at all. This enables StepCCL and GEMM to run simultaneously on a GPU without slowing down each other. This cornerstone facilitates our subsequent development of communication overlap.
Figure 13 shows an example of how StepCCL works in overlapping the allgather (AG) operation with GEMM. We start by decomposing the single GEMM and the corresponding communication into several smaller pairs. Each small communication operation starts sequentially on a communication stream, with its paired GEMM executed on the default computation stream. The communication overhead is fully hidden except for the first allgather.111If the number of allgather/GEMM is large enough, the only allgather in the critical path should have negligible overhead. But dividing a large GEMM into finer granularity sometimes could lead to overall slowdown. In practice, the number is actually configurable. After all GEMMs finish, we perform an extra layout remapping operation (usually with negligible overhead) to ensure identical results with the baseline. Figure 14 describes the details of the layout remap process. In some rare cases during backward propagation, we find the remap overhead is high due to certain model dimension. To mitigate this, we further overlap the remap with the computation of the weight gradients, so eventually we nearly get the full performance gain of the communication overlap.
Finally, although the overlap idea is also studied in many related works [42, 43, 44], we highlight the key differences of ours. Unlike prior work that fuses GEMM with TP communication into one CUDA kernel [42, 43], we choose a modular design and do not use fusion for more flexibility. For example, when TP communication is longer than GEMM, fusing them cannot fully hide the communication overhead. However, with the modular design, we are able to hide the communication with other modules without dependency (e.g., in cross-attention), which is not possible with the fused implementation. This enables broader adoption of StepCCL in many other scenarios.
8 Evaluation
@@
In this section, we first use large-scale experiments to demonstrate the overall performance improvements of DistTrain over Megatron-LM. Next, we use microbenchmarks to deep dive into DistTrain and show the effectiveness of each component in DistTrain. Finally, we provide a case study to further elucidate DistTrain’s capabilities.
@@
Setup. Our experiments are conducted on a production GPU cluster for multimodal LLM training, with each node equipped with eight NVIDIA A800 GPUs, 1TB of memory, and 96 vCPUs. GPUs within one node are interconnected by 300GB/s (bidirectional) NVLink, while nodes are connected by 4*200 Gbps RDMA network based on RoCEv2 with rail-optimized topology. The overall experiments use up to 1296 GPUs, and the microbenchmark utilizes up to 98 GPUs. We use PyTorch 2.1.2 and NVIDIA CUDA 12.2 for our evaluation.
Models. For LLM backbone, we choose the representative LLM architecture, Llama3 [22], which is widely used in both academia and industry. Table 1 lists the detailed model configurations. As for the modality, we focus on images and texts. DistTrain is also compatible with other modalities. For modality encoder and modality generator, we use ViT-Huge [45] (0.63B) and Stable-Diffusion 2.1 [46] (1B) respectively. These two models are widely used for image understanding and generation. The three LLM backbones (i.e., Llama3-7B, Llama3-13B, and Llama3-70B) are paired with ViT-Huge and Stable-Diffusion to form multimodal LLMs designated as MLLM-9B, MLLM-15B, and MLLM-72B. For large multimodal LLM (i.e., MLLM-72B), we use high image resolution (i.e., 10241024) for generation since the large LLM is able to process more context information. For small models, we use low image resolution (i.e., 512512).
@@
Datasets. For our experiments, we use the representative open-source dataset, LAION-400M. We generate training data by interleaving the image and text subsequences, forming input sequences up to 8192 tokens long. This dataset is also employed in our production multimodal LLM training. As detailed in §2.3, each training sample includes a varying number of image tokens and text tokens, which introduces data heterogeneity in multimodal LLM training.
@@
Metrics. We use the Model FLOPs Utilization (MFU) as the primary metric to evaluate DistTrain. MFU measures the percentage of GPU FLOPs that are effectively utilized during model training. We also use the training throughput (TPT.) to evaluate the training speed of DistTrain. Since DistTrain and Megatron-LM may utilize different numbers of GPUs due to varying model orchestration strategies, we also indicate the number of GPUs used in the throughput charts.
8.1 Overall Performance

### v3

@@ 5. Addressing Data Heterogeneity
To address data heterogeneity, we first introduce disaggregated data preprocessing that separates data preprocessing from training. This eliminates the interference between data processing and training and enables negligible overhead of data processing. We then introduce disaggregated data reordering to address data heterogeneity. We use intra-microbatch reordering to eliminate stragglers across DP groups, and inter-microbatch reordering to minimize pipeline bubbles and hide the training stragglers into the training pipeline as much as possible.
5.1. Disaggregated Data Preprocessing
As discussed in §2.3, preprocessing high-volume multimodal data (e.g., high resolution images) incurs substantial overhead. DistTrain disaggregates data preprocessing from training with a producer-consumer model. The producer, operating on dedicated CPU nodes, fetches data from the distributed file system and preprocesses training data asynchronously. The consumer, i.e., the main training process on dedicated GPU nodes, receives the preprocessed data for training. The producer and consumer communicate through RPC calls, and use RDMA for lower latency if available.
The disaggregation enables flexible and elastic data preprocessing and provides opportunities for data reordering without incurring additional overhead. Leveraging this opportunity, we incorporate data reordering into data preprocessing to mitigate training stragglers caused by data heterogeneity. The reordering includes two levels: intra-microbatch and inter-microbatch. We emphasize that disaggregated data preprocessing ensures that the complex reordering does not interfere with the GPU training or impose extra overhead.
5.2. Intra-microbatch Reordering
@@
Insight. To address intra-microbatch stragglers, we first identify the straggler by pinpointing the DP group with the largest training samples. As shown in Figure 6, the first DP group becomes a straggler as it contains the two largest training samples. Therefore, we reorder the training samples within the global batch by size. Specifically, as depicted in Figure 11, we reorder the training samples into the sequence , which distributes the computation more evenly.
Intra-microbatch Reordering. Leveraging this insight, we propose intra-microbatch reordering to balance the computational load within a global batch. Formally, the goal is to minimize the maximum computation time across data parallelism (DP) groups. This problem maps to the multiway number partitioning problem (Korf, 2009), which is NP-hard. Given the problem’s complexity and the large batch sizes common in production, finding an optimal solution is computationally infeasible. We therefore employ a lightweight, greedy partitioning algorithm that yields an approximation ratio of less than 4/3 (Barman and Krishnamurthy, 2020). Importantly, this reordering only occurs within the global batch, which only affects the sequence of gradient accumulation. Since gradient accumulation (i.e., summation) is a commutative operation, the intra-microbatch reordering preserves the training’s convergence semantics.
The detailed algorithm is summarized in Algorithm 1. The function receives the original training samples and DP size . This algorithm first sorts the training samples in descending order by the sample size (line 3). Then, it loops over the training samples and assigns the sample to the DP group with the current lowest computational load (line 6-8). It then returns the reordered samples (line 9-11). The algorithm has a time complexity of .
5.3. Inter-microbatch Reordering
@@
As we discussed in §2.3, data heterogeneity leads to varied computation times across microbatches within the modality encoder and generator. The straggler microbatch prolongs the training by creating large pipeline bubbles. In the context of 1F1B pipeline scheme, the overall iteration time is primarily decided by pipeline bubbles and the computation time at the first PP stage of the modality encoder, as illustrated in Figure 12. Let the PP size be and the number of microbatches be ( and in Figure 12). Typically, is larger than to reduce the proportion of time spent in the warm-up and cool-down phases. To help describe our solution, we first introduce a concept called pipeline intervals at the first PP stage. As shown in Figure 12, these intervals are typically filled with the forward pass, except for the last intervals (i.e., , , and ). Straggler microbatches in either the encoder or generator prolong these intervals or increase the unfilled area (i.e., bubble). Designing an efficient runtime algorithm to minimize pipeline bubbles presents a significant challenge.
@@
Insights. We leverage two insights to solve this problem. The first insight involves minimizing the volume of intervals that are not filled. As shown in Figure 12, the last intervals (i.e., to ) remain unfilled. These intervals become the pipeline bubbles and increase the training iteration. We observe a positive correlation between the volume of and the size of the microbatch. The size refers to the computation time of the microbatch in modality encoder and generator. For instance, is significantly larger than and since the microbatch (i.e., d) is the largest. By strategically reordering the training samples to position the smallest microbatches at the end, we are able to reduce these unfilled intervals’ volume.
@@
The second insight involves minimizing the unfilled area of the left intervals, which hides the training stragglers into the training pipeline as much as possible. The left intervals (i.e., to ) are filled with the forward pass. As shown in Figure 12, the first interval is filled with by the to forward passes. For subsequent intervals, is filled by the forward pass. However, the forward pass may not perfectly match the interval volumes, leading to unfilled areas. By evaluating the volume of , we place the microbatches, whose forward time most closely matches this volume, at the corresponding position, to minimize the unfilled area.
@@
Inter-microbatch Reordering. Based on the preceding insights, we propose a runtime algorithm for inter-microbatch reordering designed to minimize pipeline bubbles. This algorithm is specifically developed for the 1F1B pipeline schedule. We will retrofit the algorithm to VPP (i.e., interleaved 1F1B) later. Algorithm 2 summarizes the pseudo code. The function accepts the initial microbatch sequence and the PP size . The reordering process begins by scheduling the smallest microbatch first to ensure all pipeline stages are activated promptly (line 3). Subsequently, it reserves the smallest remaining microbatches for the end of the sequence to minimize unfilled intervals (lines 4 and line 12). The main loop (lines 5-11) then iterates through the remaining microbatches to fill the pipeline bubbles. In each loop iteration, it computes the interval size via the function and applies a heuristic to select microbatches that best fit this size. For the first interval, it greedily selects microbatches whose aggregate forward time most closely matches the interval size; for all subsequent bubbles, it selects the single best-fitting microbatch. This loop ensures maximal filling of the remaining intervals, which minimizes pipeline bubbles. Crucially, this reordering only alters the data sequence within the local batch of a single DP rank per training iteration. Therefore, analogous to intra-microbatch reordering, it only affects the sequence of gradient accumulation and consequently preserves the training’s convergence semantics.
@@
The functions and operate with a time complexity of . The function calculates interval size using the current order . This calculation is facilitated by a dynamic programming algorithm that utilizes a recursive formula derived from pipeline dependencies. Specifically, the start time of each microbatch depends on two factors: the completion of the preceding microbatch in the same pipeline stage and the availability of input data from the upstream microbatch. Consequently, the end time of each microbatch is determined by the maximum of these two dependencies plus its own computation time. This dynamic programming algorithm exhibits a complexity of per function invocation. The algorithm has a time complexity of .
@@
Virtual pipeline parallelism (i.e., interleaved 1F1B) also follows the one forward and one backward pipeline scheme to reduce the memory footprint. The fundamental insights of our algorithm apply to any 1F1B-based pipeline, including VPP. We adapt the algorithm by computing multiple (i.e., VPP size) intervals and filling them with the corresponding number of forward passes from a single microbatch.
6. System Implementation
We implement DistTrain with 6.3K lines of code in C++ and Python, and integrate it with Megatron-LM (Shoeybi et al., 2019). DistTrain handles failures by automatically recovering the training from the latest model checkpoint. To mitigate Tensor Parallelism (TP) overhead, we developed StepCCL, an in-house collective communication library. By leveraging the DMA engine for data transfers, StepCCL overlaps communication with computation, which frees the core Streaming Multiprocessors (SMs) to execute computation kernels (e.g., GEMM) without performance interference. A detailed description and evaluation are in §A.1.
@@ 6. System Implementation
DistTrain manager and initializer. DistTrain’s training manager, running on a dedicated CPU node, formulates the disaggregated model orchestration problem using Disciplined Convex Programming (Grant et al., 2006). It employs the CVX solver (cvx, 2024) to solve this problem within a second. The manager records the optimal resource allocation and parallelism strategy to a configuration file, which the Kubernetes controller uses to launch the training task. As for the initializer, DistTrain uses PyTorch Distributed (tor, 2024) library to initialize the communication groups. The initialization of each module in a multimodal LLM is implemented through parallelism units.
@@
Parallelism unit. As discussed in §4.1, disaggregated model orchestration is implemented via a dedicated module: parallelism unit. During distributed training initialization, DistTrain first establishes communication groups within a parallelism unit. Each GPU process possesses a global and a local rank within its unit, facilitating distributed initialization. Subsequently, DistTrain initializes a communication broker to establish PP communication between adjacent parallelism units. All inter-unit communication traffic is routed via the communication broker. The communication broker is implemented by modifying Megatron-LM’s batched send/receive operations into discrete operations. It enables flexible communication between multiple upstream and downstream GPU processes by concentrating and scattering data as needed, while preserving data order. The communication broker employs a decentralized design, residing on the GPU of either the last PP stage in an upstream unit or the first PP stage in a downstream unit. To maximize communication bandwidth, the number of brokers between two units is determined by the greatest common divisor of their respective DP sizes. Consequently, the total inter-unit bandwidth scales effectively with the training workload, preventing the communication broker from becoming a training bottleneck. Moreover, Megatron-LM’s reliance on synchronous communication compels upstream stages to pause until downstream stages fully receive data, introducing unnecessary pipeline dependencies. To mitigate this, we implement asynchronous send operations, eliminating these superfluous dependencies, and redesign the communication topology to prevent potential deadlocks.
7. Evaluation
@@
In this section, we first use large-scale experiments to evaluate the overall performance improvements of DistTrain over Megatron-LM. Next, we use an ablation study to deep dive into DistTrain and show the effectiveness of each technique. Finally, we provide a case study to further evaluate DistTrain.
Setup. Our experiments are conducted on a production cluster, with each node equipped with 8 NVIDIA Ampere GPUs. GPUs within one node are interconnected by 300GB/s (bidirectional) NVLink, while nodes are connected by 4*200 Gbps RDMA network based on RoCEv2 with rail-optimized topology. The overall experiments use up to 1296 GPUs, and the ablation study utilizes up to 96 GPUs. We use PyTorch 2.1.2 and NVIDIA CUDA 12.2. As for the modality, we focus on images and texts. We emphasize that DistTrain is also compatible with other modalities. Our evaluation focuses on images (i.e., vision). This choice is motivated by the prevalence of vision as a common non-textual modality in multimodal LLMs and the increasing prominence of capable vision-based MLLMs (e.g., Gemini (gem, 2024) and Qwen2.5-VL (Bai et al., 2025)).
Models. For the LLM backbone, we choose the representative LLM architecture, Llama3 (lla, 2024), which is widely used in both academia and industry. Table 2 lists the detailed model configurations. For encoder and generator, we use ViT-Huge (vit, 2024) (0.63B) and SD 2.1 (sd, 2024) (1B) respectively. The encoder and generator in our experiments are representative of state-of-the-art MLLM. The encoder (i.e., ViT-Huge) in our experiment aligns with the encoders of Qwen2.5-VL and Seed1.5-VL in terms of both model architecture and size. The generator (i.e., SD) is also adopted in many MLLMs (e.g., Step1X-Edit (Liu et al., 2025) and EMU (Sun et al., 2023b)) with generation capabilities. The three LLM backbones (i.e., Llama3-7B, Llama3-13B, and Llama3-70B) are paired with ViT-Huge and SD to form multimodal LLMs denoted as MLLM-9B, MLLM-15B, and MLLM-72B. For the large multimodal LLM (i.e., MLLM-72B), we use high image resolution (i.e., 10241024) for generation since the large LLM is able to process more context information. For small models, we use low image resolution (i.e., 512512).
@@
Datasets. For our experiments, we use the representative open-source dataset, LAION-400M (Schuhmann et al., 2021). We generate training data by interleaving the image and text subsequences, forming input sequences up to 8192 tokens long. This dataset is also employed in our production multimodal LLM training. As detailed in §2.3, each training sample includes a varying number of image tokens and text tokens, which introduces data heterogeneity in multimodal LLM training.
@@
Metrics. We use Model FLOPs Utilization (MFU) as the primary metric. MFU measures the percentage of GPU FLOPs that are effectively utilized during training. We also use the training throughput to evaluate the training speed of DistTrain. Since DistTrain and Megatron-LM may utilize different numbers of GPUs due to varying model orchestration strategies, we also indicate the number of GPUs used in each experiment.
Baselines. We focus on training multimodal LLMs rather than traditional multimodal models like CLIP or LiT. DistMM (Huang et al., 2024) efficiently handles traditional multimodal training but lacks support for pipelines involving encoders, LLM backbones, and generators in multimodal LLMs. Multimodal LLMs combine multimodal models and LLMs to understand and generate high-quality multimodal data auto-regressively. Megatron-LM is selected as the primary baseline due to its established role as a highly optimized, open-source framework for LLM training that also provides supports for multimodal models. For ablation study of model orchestration, we integrate DistMM’s model orchestration strategy (resource allocation by model size and FLOPs) into DistTrain, naming this baseline DistMM∗. Note that DistMM does not support multimodal LLM training natively. DistMM∗ only uses its orchestration strategy, with all other techniques from DistTrain.
7.1. Overall Performance